OpenStudy (anonymous):

I ONLY HAVE A FEW MINUTES TO POST LIKE 7 QUESTION PLEASE HELP SOMEONE!

OpenStudy (anonymous):

haha go for it

OpenStudy (anonymous):

cmon fire away

OpenStudy (anonymous):

i think she's already screwed, her online exam has probably ended at the top of the hour lol...but if not fire away like him said lol

OpenStudy (anonymous):

A CELLUAR PHONE COMPANY OFFERS A CONTACT C, IN DOLLARS, OF T MINUTES OF TELEPHONE IS GIVEN BY C=0.25(T-300) + 79.95, WHERE IT IS ASSUMED THAT T => 300 MINUTE. WHAT TIMES WILL KEEP COSTS BETWEEN $118.45 AND $164.70?

OpenStudy (anonymous):

What is this bovice?

OpenStudy (anonymous):

I think you know something about this mysterious college girl

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

i'm just answering her question

OpenStudy (anonymous):

What is all this online test. Could you please tell me

OpenStudy (anonymous):

You know this college girl, she comes to this site, and asks for answers, and only answers and no explanations

OpenStudy (anonymous):

I didn't quite understand what she is up to

OpenStudy (anonymous):

so T=454 will make C=118.45

OpenStudy (anonymous):

But I felt she is doing something she must not do

OpenStudy (anonymous):

T=639 makes C=164.70

OpenStudy (anonymous):

so the answer is 454 < T < 639

OpenStudy (anonymous):

so can i have my medal now??? (:

OpenStudy (anonymous):

why did I 301.54 and 303.39

OpenStudy (anonymous):

because you fail at basic algebra? i dunno....i thought you were in a hurry...now chop chop!!!

OpenStudy (anonymous):

TWO ANGLES ARE COMPLEMENTARY. THE SUM OF THE MEASURE OF THE FIRST ANGLE AND ONE-FOURTH THE SECOND ANGLE 77.55 DEGREE. FIND THE MEASURE OF THE ANGLES? WHAT IS THE MEASURE OF THE SMALLER ANGLE? ALSO WHAT IS MEASURE OF THE ANGLE?

myininaya (myininaya):

the one before should be 301.54<T<303.39

myininaya (myininaya):

A+B=90 A+B/4=77.55 Find A where A is your fist angle Find B where B is your second angle

OpenStudy (anonymous):

for the cost to be between $118.45 abd $164.70 telephonig time be between ? minutes and ? minutes.

OpenStudy (anonymous):

myininaya is wrong

OpenStudy (anonymous):

454 < T < 639

OpenStudy (anonymous):

118.45 = 0.25(T-300) + 79.95 subtract 79.95 from both sides.... 38.50 = 0.25(T-300) divide both sides by 0.25 154 = T - 300 add 300 to both sides T = 454

myininaya (myininaya):

C=25(T-300)+79.95 we want C to be between 118.45 and 164.7 118.45<C<164.7 118.45<25(T-300)+79.95<164.7

OpenStudy (anonymous):

I have been using this all day and I cant even get a better grade but yet you guys really know what the answer are huh funny.

OpenStudy (anonymous):

now do the same thing for the upper limit of C C = 164.70

myininaya (myininaya):

oh i see i misread the Cost sorry

OpenStudy (anonymous):

oh yeah the equation is C = 0.25(T-300) + 79.95

myininaya (myininaya):

i seen 25 and not .25 my mistake

OpenStudy (anonymous):

anyways the second part of the answer would be 164.70 = 0.25(T-300) + 79.95 84.75 = 0.25(T-300) 339 = T-300 T = 639

OpenStudy (anonymous):

w00t i got my medal :)

OpenStudy (anonymous):

I am sorry that I am stupid sorry